SS DF MS F
Among Treatments 4468.45
Error ? 10 706.87
Total 15

Step 1 of 8:

Calculate the sum of squares of experimental error. Please round your answer to two decimal places.

Step 2 of 8:

Calculate the degrees of freedom among treatments.

Step 3 of 8:

Calculate the mean square among treatments. Please round your answer to two decimal places.

step 4 of 8

Calculate the F-value. Please round your answer to two decimal places.

step 5 of 8

What is the sum of squares of sample means about the grand mean? Please round your answer to two decimal places.

Step 6 of 8:

What is the variation of the individual measurements about their respective means? Please round your answer to two decimal places.

Step 7 of 8:

What is the critical value of F at the 0.05 level? Please round your answer to four decimal places, if necessary.

Step 8 of 8:

Is F significant at 0.05?
Answer #1

Step 1 of 8:

Calculate the sum of squares of experimental error. Please round your answer to two decimal places.

sum of squares of experimental error = DF(error) * MS(Error) = 10 * 706.87 = 7068.70

Step 2 of 8:

Calculate the degrees of freedom among treatments.

Degree of freedom among treatment = Df(total) - Df(error) = 15 - 10 = 5

Step 3 of 8:

Calculate the mean square among treatments. Please round your answer to two decimal places.

Mean square among treatments = SSE/ dF(Treatment) = 4468.45/ 5 = 893.69

step 4 of 8

Calculate the F-value. Please round your answer to two decimal places.

F = 893.69/ 707.87 = 1.2643

step 5 of 8

What is the sum of squares of sample means about the grand mean? Please round your answer to two decimal places.

Here Sum of squares of sample means about the grand mean = 4468.45

Step 6 of 8:

What is the variation of the individual measurements about their respective means? Please round your answer to two decimal places.

the variation of the individual measurements about their respective means = 7068.7

Step 7 of 8:

What is the critical value of F at the 0.05 level? Please round your answer to four decimal places, if necessary.

F (critical) = F(0.05,5,10) = 3.3258

Step 8 of 8:

Is F significant at 0.05?

Here F < F(critical) so no, F is not significant at 0.05 level.